OpenStudy (anonymous):

Help please i will give you medals!!! I need help asap please!!!! What is the value is a discontinuity of x^2+2x+3 over x^2-x-12 x=-1 x=-2 x=-3 x=-4

jimthompson5910 (jim_thompson5910):

Factor x^2+2x+3 to get (x+1)(x+2) Factor x^2-x-12 to get (x-4)(x+3) Nothing simplifies/cancels, so we focus on the denominator now. The denominator is factored to be (x-4)(x+3). Set it equal to zero and solve for x (x-4)(x+3) = 0 x-4=0 or x+3=0 x=4 or x=-3 Those last two values, x=4 or x=-3, make the denominator zero...but remember: you CANNOT divide by zero. So those values cause a discontinuity

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

so is it -3?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

can you help me with another one?

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

Find the horizontal or oblique asymptote of f(x) = negative 3 x squared plus 7 x plus 1, all over x minus 2 y = 2 y = -3 y = 3x + 7 y = -3x + 1

jimthompson5910 (jim_thompson5910):

Are you familiar with polynomial long division?

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

for a rational, a "discontinuity" will be at the point where its asymptotes are as pointed out by jim_thompson5910 , you can get the horizontal asymptotes by getting the "zeros" of the denominator, that is \(\bf x^2-x-12 = 0\) at those points, the denominator will be zero, thus the fraction "undefined", thus will yield a "discontinuity"

OpenStudy (whpalmer4):

@jim_thompson5910 and @jdoe0001 have really said it all...

OpenStudy (triciaal):

please note there was an error in the first line x^2 + 2x + 3 is not (x + 1)(x+2) it did not affect the denominator